Suppose the government decides to issue a new savings bond that is guaranteed to double in value if you hold it for 20 years. Assume you purchase a bond that costs $75. a. What is the exact rate of return you would earn if you held the bond for 20 years until it doubled in value? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g.,32.16.) year, how much would the bond be worth in 2027? (Do not round intermediate calculations and round your answer to 2 decimal places,e.g, 32.16.) c. In 2027, instead of cashing in the bond for its then current value, you decide to hold the bond until it doubles in face value in 2037. What annual rate of return will you earr over the last 10 years? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g,32.16.) a. Rate of return b. Value of bond C. Rate of return
Answer #1

Answer a.

Price in 2017 = $75
Price in 2037 = $150
Time Period = 20 years

Rate of Return = (Price in 2037 / Price in 2017)^(1/20) - 1
Rate of Return = ($150 / $75)^(1/20) - 1
Rate of Return = 2^(1/20) - 1
Rate of Return = 1.0353 - 1
Rate of Return = 0.0353 or 3.53%

Answer b.

Rate of Return during first 10 years = 0.13%

Price in 2027 = Price in 2017 * (1 + Rate of Return)^10
Price in 2027 = $75 * 1.0013^10
Price in 2027 = $75.98

Answer c.

Price in 2027 = $75.98
Price in 2037 = $150
Time Period = 10 years

Rate of Return = (Price in 2037 / Price in 2027)^(1/10) - 1
Rate of Return = ($150 / $75.98)^(1/10) - 1
Rate of Return = 1.974204^(1/10) - 1
Rate of Return = 1.0704 - 1
Rate of Return = 0.0704 or 7.04%